# Simultaneous analysis of approximate leave-one-out cross-validation and mean-field inference

### v1

@@ 1.4. Examples of matrices $\hat{A}$
For concreteness, let us provide the expression of $\hat{A}$ in a few examples of interest. Assuming that $L_{y_{i}}(\cdot)$ is twice-differentiable, and let $D\in\mathbb{R}^{n\times n}$ be the diagonal matrix (1.9) with $D_{ii}=L_{y_{i}}^{\prime\prime}(x_{i}^{T}\hat{b})$.
@@
If $R$ is twice differentiable with Hessian $\nabla^{2}R(b)$ have smaller eigenvalue bounded from below by $\nu n$ for some $\nu\geq 0$, for every $b\in\mathbb{R}^{p}$. Here, $\hat{A}=(X^{T}DX+\nabla^{2}R(\hat{b}))^{-1}.$
@@
almost surely where $\hat{S}=\{j\in[p]:\hat{b}_{j}\neq 0\}$ is the support of $\hat{b}$. Above $\hat{X}_{\hat{S}}$ is the submatrix of $X$ with columns indexed in $\hat{S}$ and $I_{\hat{S},\hat{S}}$ is the identity matrix of size $|\hat{S}|$.
@@
If $\lambda_{1},...,\lambda_{k}>0\nu\geq 0$, a partition $(G_{1},...,G_{K})$ of $[p]$ is given and $R(b)=\sum_{k=1}^{K}\lambda_{k}\|b_{G_{k}}\|_{2}+n\nu\|b\|_{2}^{2}/2$ then with $\hat{T}=\{k\in[K]:\|\hat{b}_{G_{k}}\|_{2}>0\}$ and $\hat{S}=\cup_{k\in\hat{T}}G_{K}$, almost surely
@@
These formulae are obtained by differentiating the KKT conditions. For instance, in the case of twice-differentiable $R$, differentiating $\sum_{l=1}^{n}x_{l}L_{y_{l}}^{\prime}(x_{l}^{T}\hat{b})+\nabla R(\hat{b})=0$ with respect to $x_{ij}$ gives $e_{j}L_{y_{i}}^{\prime}(x_{i}^{T}\hat{b})+(\nabla^{2}R(\hat{b})+X^{T}DX)\frac{\partial}{\partial x_{ij}}\hat{b}+X^{T}De_{i}\hat{b}_{j}=0$. The derivative formula in Proposition 1.1 thus holds with $\hat{A}$ given by Example 1.1. More care is needed in the case of Examples 1.2 and 1.3 as the penalty is non-differentiable. In such non-differentiable case, the above formulae are obtained by first showing that almost surely the active set $\hat{S}$ is locally constant; once this is established, the KKT conditions can be differentiated as in the twice-differentiable case. We refer to [9, Section 4] for such arguments applied to the lasso and group-lasso. In Proposition 1.1 the matrix $\hat{A}$ is shown to exist, however no general closed-form expression for $\hat{A}$ is available for general regularizer $R$ at this point. The above cases are the most common though, and the formulae for $\hat{A}$ in these cases are explicit and given above.
@@
then the expression of $\hat{A}$ in the derivative formula of Proposition 1.1 and the expression of $\hat{A}$ in the ALO-CV construction (1.7) coincide.
In existing analysis of ALO-CV for non-differentiable regularizer [1] where the penalty is the elastic-net, the major difficulty comes from controlling the difference between the support $\hat{S}$ of $\hat{b}$ defined in Example 1.2 and the support $S^{i}$ of the leave-one-out vector $b^{i}$ in (1.4). Auddy et al. [1] manages to achieve such control between the supports $S^{i}$ and $\hat{S}$ for the square loss, elastic-net penalty and isotropic Gaussian designs, and as a consequence obtains the consistency of ALO-CV in this case. This paper develops an alternative argument to derive the consistency of ALO-CV for non-differentiable regularizer, by relating ALO-CV directly to mean-field approximations such as (1.13), without trying to control the difference between the supports $\hat{S}$ and $S^{i}$ or other discrepancies between $\hat{b}$ and $b^{i}$.
@@ 1.5. Reconciling ALO-CV and mean-field inference
Typical proportional asymptotics results such as (1.8) (e.g., Theorems 1 and 2 in [22]) also suggest that the weight in front of $L_{y_{i}}^{\prime}(x_{i}^{T}\hat{b})$ should be independent of $i$ to obtain the mean-field limit of the generalization error [22, Theorem 1, eq. (2.5)].
@@
see [5, Section 5] in linear regression and [3] in single-index models. Without the two sums in the numerator and the denominator, the rightmost fraction would be the same as the $W_{i}$ weights from ALO-CV. The approximation (1.18) in [5, 3] is proved for nondifferentiable regularizers, so (1.18) provides a little hope towards reconciling ALO-CV with (1.17) under no smoothness assumption on the regularizer $R(\cdot)$ in (1.1). However, the proofs of (1.18) in [3, 5] do leverage the two summations in the numerator and denominator of (1.18). The gist of the present paper is a novel argument, different from the leave-one-out bounds in [34, 1] or the Stein formulae argument to prove (1.18) in [5, 3], to obtain $W_{i}\approx\operatorname{tr}[\Sigma\hat{A}]$ for non-differentiable regularizers.
@@ 1.6. Notation
[…]
@@ 2. Square loss and rotational invariance
On the other hand, [5, Section 5] proves that under a $n\mu$-strong-convexity assumption on $R$ and (1.3),
@@
with probability at least $1-c/n$ for some absolute constant $c>0$. Furthermore, if $R$ in (2.1) is $(\mu n,I_{p})$-strongly convex then in the same event
@@
can be used interchangeably in results of the form (1.17) up to a small error term of order $\sqrt{\log(n)/n}$. The proof given below relies mostly on the rotational invariance of the square loss and the Gaussian distribution. Smoothness of the regularizer is irrelevant. Inequality (2.3) only requires convexity of the regularizer, while for (2.4) we additionally assume strong-convexity as this is a simple condition that lets us control $1/(1-\operatorname{tr}[H]/n)$. Controlling $1/(1-\operatorname{tr}[H]/n)$ without strong-convexity is possible, e.g., [14], however this requires more involved arguments and different assumptions such as small-enough sparsity of $\beta^{*}$ [14, 4].
Let us explain with an example how (2.3) can be used with existing results in mean-field asymptotics to draw conclusions about ALO-CV cross-validation. By Theorems 7 and 8 in [14], for any 1-Lipschitz function $\phi:\mathbb{R}^{n}\to\mathbb{R}$, if the sparsity of $\beta^{*}$ is small enough (in the sense of the phase transition studied in [14]), the Lasso $\hat{b}$ satisfies the convergence in probability
@@
and $1+W_{i}=1/(1-H_{ii})$. Theorem 8 in [14] shows that $\operatorname{tr}[H]/n=\|\hat{b}\|_{0}/n$ is bounded away from 1 in the sense that $01-\operatorname{tr}[H]/n>c_{0}>$ 0 with high-probability, so that by (2.3) the scalar $H_{ii}$ is also bounded away from 1 uniformly in $i\in[n]$. Since $\varphi$ is Lipschitz, (2.3) gives
@@
$H$ $=\tfrac{\partial}{\partial y}(X\hat{b})$ by definition of $H$
$=Q\tfrac{\partial}{\partial y}(G\hat{b})$ by linearity
@@ 3. Beyond rotational invariance: probabilistic results
The proof is given in Appendix A In other words, $\operatorname{Rem}_{i}$ concentrates exponentially fast (although with $t^{3/2}$ inside the exponential) around its conditional expectation $\mathbb{E}[\operatorname{Rem}_{i}\mid E]$. Using a union bound over $i\in[n]$ taking $t^{2/3}=C\log(n^{2})$ in (3.3), we get with probability $1-C/n$ that
@@
Multiplying by $-D_{ii}$ and adding 1, inequality (3.5) is equivalent to
@@
Restriction to a support or a subset
@@
for some orthogonal projection $W\in\mathbb{R}^{n\times n}$. If $S\subset[p]$ is a fixed subset of indices and $W=\sum_{j\in S}e_{j}e_{j}^{T}$, the above limit is given by
$A_{S,S}=(X_{S}^{T}DX_{S}+(\Sigma^{1/2}H\Sigma^{1/2})_{S,S})^{-1},\qquad A_{jk}=0\text{ if }j\notin S\text{ or }k\notin S.$ (3.8)
@@
With $A_{t}=(X^{T}DX+\Sigma^{1/2}H\Sigma^{1/2}+tW)^{-1}$ for $t\geq 0$, we have the monotone limits $\operatorname{tr}[\Sigma^{1/2}A_{t}\Sigma^{1/2}]\downarrow\operatorname{tr}[\Sigma^{1/2}A\Sigma^{1/2}]$ and $x_{i}^{T}A_{t}x_{i}\downarrow\operatorname{tr}[A]$ as $t\uparrow+\infty$. Applying Theorem 3.1 to $A_{t}$ and taking $t\to+\infty$ using the dominated convergence theorem gives Corollary 3.2.
@@
The notation and setting used so far in Section 3 is seemingly detached from the notation in Examples 1.2 and 1.3 or Proposition 1.1. In the result Theorem 3.1, the matrices $D$ and $H$ are deterministic, and the support $S$ in (3.8) is also deterministic for Corollary 3.2 to hold with the deterministic projection $W=\sum_{j\in S}e_{j}e_{j}^{T}$.
@@
which is random because $\nabla^{2}R(\hat{b})$ is random through $\hat{b}$, and $D=\operatorname{diag}(\{L_{y_{i}}^{\prime\prime}(x_{i}^{T}\hat{b}),i\in[n]\})$ is also random through $L_{y_{i}}^{\prime\prime}(x_{i}^{T}\hat{b}),i\in[n]$. In Examples 1.2 and 1.3, the support $\hat{S}=\{j\in[p]:\hat{b}_{j}\neq 0\}$ is random as well, and it is unclear at this point why Theorem 3.1 and Corollary 3.2 with a deterministic $S\subset[p]$ in (3.8) would be useful at all. This will be resolved by considering particular events of the form (3.1), such that in this event $D,\hat{b},\nabla^{2}R(\hat{b})$, and in the case of Examples 1.2 and 1.3 the support $\hat{S}$, are all fully determined (i.e., they are conditionally conditionally on an event of the form (3.1)). Such events have been studied in [12, 13] and are described in the following sections.
@@ 4. Conditioning example I: Robust linear regression
The key intuition behind the argument of Proposition 2.1 is that in the square loss case, if $Q\Lambda P^{T}$ is the SVD of the design matrix $[X\mid\varepsilon]\in\mathbb{R}^{n\times(p+1)}$, then $\hat{b}$ and $\operatorname{tr}[H]$ are independent of the left singular vectors $Q$. We can then argue conditionally on $(\Lambda,P)$ with respect to the randomness of the randomness of the independent matrix $Q$.
@@
are satisfied. Hence in this event $\hat{b}=\bar{h}+\beta^{*}$, $X(\hat{b}-\beta^{*})=\bar{u}$ and $\bar{w}$ is an element of $\partial R(\hat{b})$ that satisfies the KKT conditions for $\hat{b}$. The second key observation made in [12, 13] is that the event (4.4) is made of linear constraints on the Gaussian matrix $X$. By properties of the multivariate normal distribution, conditionally on (4.4), the entries of $X$ are again jointly normal.
The event (4.4) lets us essentially condition on $\hat{b},X(\hat{b}-\beta^{*})$ and $X^{T}\rho^{\prime}(y-X\hat{b})$, while maintaining that the entries of $X$ are jointly normal (although with different mean and covariance than the original random matrix).
@@
Let $\mu,\delta>0$ be constants. Assume the proportional regime (1.3). Consider the robust regression setting (4.1)-(4.1) with $\varepsilon$ independent of $X$ and loss $\rho$ differentiable with $\rho^{\prime}$ 1-Lipschitz. Assume that $X$ has iid $N(0,\Sigma)$ rows with $\Sigma$ invertible. Consider one of the regularizer in Examples 1.1, 1.2 and 1.3 and assume $\nu\|\Sigma\|_{op}\geq\mu$ so that the strong-convexity (1.10) holds. For the explicit matrix $\hat{A}$ given in Examples 1.1, 1.2 and 1.3 we have
@@
Since $\varepsilon$ is independent of $X$, we argue conditionally on $\varepsilon$ and consider $\varepsilon$ fixed in what follows. Consider deterministic $\bar{h}\in\mathbb{R}^{p},\bar{u}\in\mathbb{R}^{n},\bar{v}\in\mathbb{R}^{n},\bar{w}\in\mathbb{R}^{p}$ such that (4.3) holds and the event (4.4). In this event, $\hat{b},D$ and $\hat{S}$ are deterministic, Choose $P,Q$ in Theorem 3.1 be projections of rank $p-1$ and $n-1$ such that (3.1) is the same as (4.4). In the case of Examples 1.2 and 1.3, apply Corollary 3.2 with $W=\sum_{j\in\hat{S}}e_{j}e_{j}^{T}$, where $\hat{S}$ is deterministic thanks to conditioning. The bound follow from (3.3), (3.4) combined with Markov’s inequality, and (3.6) combined with $\mathbb{P}(x_{i}^{T}\Sigma^{-1}x_{i}>4p)\leq\exp(-p)$ [21, Lemma 1] and the union bound over $i\in[n]$. Once the result is proved conditionally on (3.1), it remains to integrate over $(\bar{h},\bar{u},\bar{v},\bar{w},\varepsilon)$ with respect to the probability distribution of $\hat{b}-\beta^{*},X(\hat{b}-\beta^{*}),\rho(y-X\hat{b}),X^{T}\rho^{\prime}(y-X\hat{b}),\varepsilon)$.
@@
(4.5)$\Bigl{(}\frac{1}{n}\sum_{i=1}^{n}1+\Bigl{(}x_{i}^{T}\hat{b}+W_{i}L_{y_{i}}(x_{i}^{T}\hat{b})\Bigr{)}^{2}+\Bigl{(}x_{i}^{T}\hat{b}+\operatorname{tr}[\Sigma\hat{A}]L_{y_{i}}(x_{i}^{T}\hat{b})\Bigr{)}^{2}\Bigr{)}^{1/2}.$
By $\operatorname{tr}[\Sigma\hat{A}]\leq\frac{p}{n\mu}$, (4.5) to bound $W_{i}$, $|L_{y_{i}}^{\prime}(x_{i}^{T}\hat{b})|\leq 1$ and (1.15) to bound $\frac{1}{n}\|X\hat{b}\|_{2}^{2}$ we obtain (4.6). ∎

### v2

@@ 1.4. Examples of matrices $\hat{A}$
[…]
@@
Assume $R$ is twice differentiable with Hessian $\nabla^{2}R(b)$ having smaller eigenvalue bounded from below by $\nu n$ for some $\nu\geq 0$, for every $b\in\mathbb{R}^{p}$. Then, $\hat{A}=(X^{T}DX+\nabla^{2}R(\hat{b}))^{-1}.$
@@
[…]
@@
If $\lambda_{1},...,\lambda_{k}>0$, $\nu\geq 0$, a partition $(G_{1},...,G_{K})$ of $[p]$ is given and $R(b)=\sum_{k=1}^{K}\lambda_{k}\|b_{G_{k}}\|_{2}+n\nu\|b\|_{2}^{2}/2$ then with $\hat{T}=\{k\in[K]:\|\hat{b}_{G_{k}}\|_{2}>0\}$ and $\hat{S}=\cup_{k\in\hat{T}}G_{k}$, almost surely
@@
These formulae are obtained by differentiating the KKT conditions. For instance, in the case of twice-differentiable $R$, differentiating $\sum_{l=1}^{n}x_{l}L_{y_{l}}^{\prime}(x_{l}^{T}\hat{b})+\nabla R(\hat{b})=0$ with respect to $x_{ij}$ gives $e_{j}L_{y_{i}}^{\prime}(x_{i}^{T}\hat{b})+(\nabla^{2}R(\hat{b})+X^{T}DX)\frac{\partial}{\partial x_{ij}}\hat{b}+X^{T}De_{i}\hat{b}_{j}=0$. The derivative formula in Proposition 1.1 thus holds with $\hat{A}$ given by Example 1.1. More care is needed in the case of Examples 1.2 and 1.3 as the penalty is non-differentiable. In such non-differentiable cases, the above formulae are obtained by first showing that almost surely the active set $\hat{S}$ is locally constant; once this is established, the KKT conditions can be differentiated as in the twice-differentiable case. We refer to [9, Section 4] for such arguments applied to the lasso and group-lasso. In Proposition 1.1 the matrix $\hat{A}$ is shown to exist, however no general closed-form expression for $\hat{A}$ is available for general regularizer $R$ at this point. The above cases are the most common though, and the formulae for $\hat{A}$ in these cases are explicit and given above.
@@
the expression of $\hat{A}$ in the derivative formula of Proposition 1.1 and the expression of $\hat{A}$ in the ALO-CV construction (1.7) coincide.
In existing analysis of ALO-CV for non-differentiable regularizer [1] where the penalty is the elastic-net, the major difficulty comes from controlling the difference between the support $\hat{S}$ of $\hat{b}$ defined in Example 1.2 and the support $S^{i}$ of the leave-one-out vector $b^{i}$ in (1.4). Auddy et al. [1] manage to achieve such control between the supports $S^{i}$ and $\hat{S}$ for the square loss, elastic-net penalty and isotropic Gaussian designs, and as a consequence obtains the consistency of ALO-CV in this case. This paper develops an alternative argument to derive the consistency of ALO-CV for non-differentiable regularizer, by relating ALO-CV directly to mean-field approximations such as (1.13), without trying to control the difference between the supports $\hat{S}$ and $S^{i}$ or other discrepancies between $\hat{b}$ and $b^{i}$.
@@ 1.5. Reconciling ALO-CV and mean-field inference
[…]
@@
see [5, Section 5] in linear regression and [4] in single-index models. Without the two sums in the numerator and the denominator, the rightmost fraction would be the same as the $W_{i}$ weights from ALO-CV. The approximation (1.18) in [5, 4] is proved for nondifferentiable regularizers, so (1.18) provides a little hope towards reconciling ALO-CV with (1.17) under no smoothness assumption on the regularizer $R(\cdot)$ in (1.1). However, the proofs of (1.18) in [4, 5] do leverage the two summations in the numerator and denominator of (1.18). The main technical contribution of the present paper is a novel argument, different from the leave-one-out bounds in [34, 1] or the Stein formulae argument to prove (1.18) in [5, 4], to obtain $W_{i}\approx\operatorname{tr}[\Sigma\hat{A}$] for non-differentiable regularizers.
@@ 1.6. Notation
The norms $\|\cdot\|_{2},\|\cdot\|_{1}$ are the Euclidean and $\ell_{1}$ norms of vectors. If none is specified, $\|\cdot\|$ denotes the Euclidean norm. For matrices, $\|\cdot\|_{op}$ is the operator norm (largest singular value) and $\|\cdot\|_{F}$ is the Frobenius norm. We use the set notation $n]=\{1,...,n\}$ and similarly for $[p]$.
@@ 2. Square loss and rotational invariance
On the other hand, [5, Section 5] proves that under a $n\mu$-strong convexity assumption on $R$ and (1.3),
@@
with probability at least 1c/n1-c/n for some absolute constant $c>0$. Furthermore, if $R$ in (2.1) is $(\mu n,I_{p})$-strongly convex then in the same event
@@
can be used interchangeably in results of the form (1.17) up to a small error term of order $\sqrt{\log(n)/n}$. The proof given below relies mostly on the rotational invariance of the square loss and the Gaussian distribution. Smoothness of the regularizer is irrelevant. Inequality (2.3) only requires convexity of the regularizer, while for (2.4) we additionally assume strong convexity as this is a simple condition that lets us control $1/(1-\operatorname{tr}[H]/n)$. Controlling $1/(1-\operatorname{tr}[H]/n)$ without strong convexity is possible, e.g., [14], however this requires more involved arguments and different assumptions such as small-enough sparsity of $\beta^{*}$ [14, 3].
Let us explain with an example how (2.3) can be used with existing results in mean-field asymptotics to draw conclusions about ALO-CV. By Theorems 7 and 8 in [14], for any 1-Lipschitz function $\phi:\mathbb{R}^{n}\to\mathbb{R}$, if the sparsity of $\beta^{*}$ is small enough (in the sense of the phase transition studied in [14]), the Lasso $\hat{b}$ satisfies the convergence in probability
@@
and $1+W_{i}=1/(1-H_{ii})$. Theorem 8 in [14] shows that $\operatorname{tr}[H]/n=\|\hat{b}\|_{0}/n$ is bounded away from 1 in the sense that $1-\operatorname{tr}[H]/n>c_{0}>0$ with high-probability, so that by (2.3) the scalar $H_{ii}$ is also bounded away from 1 uniformly in $i\in[n]$. Since $\varphi$ is Lipschitz, (2.3) gives
@@
[…]
@@ 3. Beyond rotational invariance: probabilistic results
The proof is given in Appendix A In other words, $\operatorname{Rem}_{i}$ concentrates exponentially fast (although with $t^{3/2}$ inside the exponential) around its conditional expectation $\mathbb{E}[\operatorname{Rem}_{i}\mid E]$. Using a union bound over $i\in[n]$ taking $t^{2/3}=C\log(n^{2})$ in (3.3), we get with probability 1C/n1-C/n that
@@
Multiplying by $D_{ii}$ and adding 1, inequality (3.5) is equivalent to
@@
Restriction to a support or a subspace
@@
for some orthogonal projection $W\in\mathbb{R}^{n\times n}$. If $S\subset[p]$ is a fixed subset of indices and $W=\sum_{j\in S^{c}}e_{j}e_{j}^{T}$, the above limit is given by
$A_{S,S}=(X_{S}^{T}DX_{S}+(\Sigma^{1/2}H\Sigma^{1/2})_{S,S})^{-1},\qquad A_{jk}=0\text{ if }j\notin S\text{ or }k\notin S,$ (3.8)
for instance by applying the block inversion formula before taking $t\to+\infty$.
@@
With $A_{t}=(X^{T}DX+\Sigma^{1/2}H\Sigma^{1/2}+tW)^{-1}$ for $t\geq 0$, we have the monotone limits $\operatorname{tr}[\Sigma^{1/2}A_{t}\Sigma^{1/2}]\downarrow\operatorname{tr}[\Sigma^{1/2}A\Sigma^{1/2}]$ and $x_{i}^{T}A_{t}x_{i}\downarrow x_{i}^{T}Ax_{i}$ as $t\uparrow+\infty$. Applying Theorem 3.1 to $A_{t}$ and taking $t\to+\infty$ using the dominated convergence theorem gives Corollary 3.2.
@@
The notation and setting used so far in Section 3 is seemingly detached from the notation in Examples 1.2 and 1.3 or Proposition 1.1. In the result Theorem 3.1, the matrices $D$ and $H$ are deterministic, and the support $S$ in (3.8) is also deterministic for Corollary 3.2 to hold with the deterministic projection $W=\sum_{j\in S^{c}}e_{j}e_{j}^{T}$.
@@
[…]
@@ 4. Conditioning example I: Robust linear regression
The key intuition behind the argument of Proposition 2.1 is that in the square loss case, if $Q\Lambda P^{T}$ is the SVD of the design matrix $X\mid\varepsilon]\in\mathbb{R}^{n\times(p+1)}$, then $\hat{b}$ and $\operatorname{tr}[H]$ are independent of the left singular vectors $Q$. We can then argue conditionally on $(\Lambda,P)$ with respect to the randomness of the randomness of the independent matrix $Q$.
@@
[…]
@@
Let $\mu,\delta>0$ be constants. Assume the proportional regime (1.3). Consider the robust regression setting (4.1)-(4.1) with $\varepsilon$ independent of $X$ and loss $\rho$ differentiable with $\rho^{\prime}$ 1-Lipschitz. Assume that $X$ has iid $N(0,\Sigma)$ rows with $\Sigma$ invertible. Consider one of the regularizer in Examples 1.1, 1.2 and 1.3 and assume $\nu\|\Sigma\|_{op}\geq\mu$ so that the strong convexity (1.10) holds. For the explicit matrix $\hat{A}$ given in Examples 1.1, 1.2 and 1.3 we have
@@
Since $\varepsilon$ is independent of $X$, we argue conditionally on $\varepsilon$ and consider $\varepsilon$ fixed in what follows. Consider deterministic $\bar{h}\in\mathbb{R}^{p},\bar{u}\in\mathbb{R}^{n},\bar{v}\in\mathbb{R}^{n},\bar{w}\in\mathbb{R}^{p}$ such that (4.3) holds, and consider the event (4.4). Conditionally on this event, $\hat{b},D$ and $\hat{S}$ are deterministic. Choose $P,Q$ in Theorem 3.1 to be projections of rank p1p-1 and n1n-1 such that the event (3.1) is the same as (4.4). In the case of Examples 1.2 and 1.3, apply Corollary 3.2 with $W=\sum_{j\in\hat{S}^{c}}e_{j}e_{j}^{T}$, where $\hat{S}$ is conditionally deterministic given the event (3.1). The bound then follows from (3.3), (3.4) combined with Markov’s inequality, and (3.6) combined with $\mathbb{P}(x_{i}^{T}\Sigma^{-1}x_{i}>4p)\leq\exp(-p)$ [21, Lemma 1] and the union bound over $i\in[n]$. Once the result is proved conditionally on (3.1), it remains to integrate over $(\bar{h},\bar{u},\bar{v},\bar{w},\varepsilon)$ with respect to the probability distribution of $(\hat{b}-\beta^{*},X(\hat{b}-\beta^{*}),\rho(y-X\hat{b}),X^{T}\rho^{\prime}(y-X\hat{b}),\varepsilon)$.
@@
[…]
By $\operatorname{tr}[\Sigma\hat{A}]\leq\frac{p}{n\mu}$, using (4.5) to bound $W_{i}$, inequality $|L_{y_{i}}^{\prime}(x_{i}^{T}\hat{b})|\leq 1$ by assumption on the loss and (1.15) to bound $\frac{1}{n}\|X\hat{b}\|_{2}^{2}$ we obtain (4.6). ∎